# Stabilizing Non-Abelian Topological Order against Heralded Noise via Local Lindbladian Dynamics

### v1

@@ IV.1 $D_{4}$ Topological Order: Quasi-stabilizer model and anyons
Figure 6: Stabilizers for the $D_{4}$ model: qubits are located on edges of three interpenetrating honeycomb lattices, depicted in red, blue, and green. We use PBC in both vertical and horizontal directions, as defined in Eq. (5), for each of these lattices. A vertex operator $B_{v_{c}}$ on sublattice $1$ of each lattice color is shown in the upper left; the vertex operators on sublattice $2$ are products of $Z$ operators on the remaining three edges at each vertex. A plaquette operator $A_{p_{c}}$ for each color is shown on the bottom left. Pauli $X$ operators act on the shaded $c$ colored edges; bold black lines indicate the two qubit control-Z gates ($CZ_{j,k}=\frac{1}{2}(1+Z_{j}+Z_{k}-Z_{j}Z_{k})$) acting on pairs of qubits inside the plaquette. The box in the upper right shows plaquette operators that fail to commute, and instead satisfy Eq. (22) with $c=\text{green},c^{\prime}=\text{blue},c^{\prime\prime}=\text{red}$.
We begin with a brief overview of the properties of $D_{4}$ topological order, as well as the operators that can be used to stabilize $D_{4}$ topologically ordered ground states on the lattice in a conventional qubit architecture. $D_{4}$ TO describes the phase of matter obtained by considering a discrete gauge theory with a gauge group $D_{4}$ (i.e. the dihedral group with 8 elements) in $d=2$ spatial dimensions. The point-like excitations, or anyons, in this theory can be understood in terms of their (electric) charge and (magnetic) flux. The pure flux excitations are associated with the four non-trivial conjugacy classes of $D_{4}$, while pure charges are associated with the four non-trivial irreducible representations (irreps) of $D_{4}$. The remaining excitations, known as dyons, are labeled by a choice of conjugacy class together with a non-trivial representation of its centralizer. For $D_{4}$, this leads to a total of twenty-one non-trivial anyons (or twenty-two total anyons, including the identity). A more comprehensive description of this anyon theory is provided in Appendix B.
Since $D_{4}$ is a non-Abelian group, some of these anyons are non-Abelian: when a pair of such anyons is brought together (or fused), the resulting anyon type is not uniquely determined. For example, $D_{4}$ has one 2-dimensional irrep; two charges of this type can be combined in multiple ways, yielding different charge types (i.e. multiple irreps), in a manner analogous to the addition of angular momentum for irreps of SU(2).
@@
In the remainder of this Section we review a specific model Hamiltonian, introduced in Ref. [104], that realizes the $\mathcal{D}(D_{4})$ topological phase. This model has two advantages: first, the local Hilbert space is composed of qubits, rendering it simpler to realize on current quantum hardware [64]; and second, the Hamiltonian can be written as a sum of local “check” operators, each of which individually take eigenvalue $1$ in the ground state subspace despite not commuting in the full Hilbert space.
The model is defined on three inter-penetrating honeycomb lattices with a qubit placed on each edge, as shown in Fig. 6. For convenience, the three honeycomb lattices are shown in different colors (red, green, and blue in the Figure). With each color, we associate two types of check operators which stabilize the ground states. First, the vertex check centered at vertex $v_{c}$ is defined as
@@
Note that the plaquette check operators defined above are non-Pauli “quasi-stabilizers”. Specifically, unlike standard stabilizer operators on an $N$ qubit Hilbert space, these check operators are not elements of the $N$ qubit Pauli group, and do not commute with each other as operators. However, they do commute within the ground-state subspace (see below for details), and stabilize the $D_{4}$ ground states. Thus in the interest of brevity (and in an abuse of notation), we will often refer to $A_{p}$ and $B_{v}$ operators simply as stabilizers and will emphasize the non-trivial consequences of their “quasi-stabilizer” nature whenever necessary.
@@
$[B_{v_{c}},B_{v^{\prime}_{c^{\prime}}}]=[B_{v_{c}},A_{p_{c^{\prime}}}]=0\,\,\forall\,p,v,v^{\prime},c,c^{\prime}\,.$ (21)
@@
$\begin{split}A_{p_{r}}X_{j_{b}}|\psi_{D_{4}}\rangle&=X_{j_{b}}Z_{k_{g}}Z_{k^{\prime}_{g}}A_{p_{r}}|\psi_{D_{4}}\rangle\\ &=X_{j_{b}}Z_{k_{g}}Z_{k^{\prime}_{g}}|\psi_{D_{4}}\rangle\,,\end{split}$ (25)
where $A_{p_{r}}$ is red plaquette that encloses the blue edge $j_{b}$ and $k_{g},k^{\prime}_{g}\in\ ^{*}\mathcal{I}(p_{r})$ denote the green edges inside this red plaquette that are connected by $CZ$ gates to the edge $j_{b}$. Thus, in configurations with $Z_{k_{g}}Z_{k^{\prime}_{g}}=+1$, measuring $A_{p_{r}}$ after the $X_{j_{b}}$ error acts on the ground state $|\psi_{D_{4}}\rangle$ will find no plaquette error, while in configurations with $Z_{k_{g}}Z_{k^{\prime}_{g}}=-1$ a plaquette error will be found. Since both configurations occur in the ground state, $X_{j_{b}}$ leaves this red plaquette in a superposition of having and not having a plaquette defect. Similarly, acting with a longer string of $X$ operators on blue edges will result in a pair of vertex violations at the string’s endpoints, together with a superposition of violated and unviolated plaquette stabilizers along the string’s length, as depicted by the highlighted plaquettes in Fig. 8(a).
@@
Here, $C_{b}$ is an open string along the blue honeycomb lattice that connects vertices $u$ and $v$, and $\ell_{b}\in C_{b}$ denotes links along this path (see Fig. 7(a)). The set $C_{rg}$ consists of those links on the red and green lattices which are incident on vertices along $C_{b}$ and are also either enclosed by or are on the boundary of two plaquettes that enclose vertices in $C_{b}$ [^5] Let $\{V_{c}\}$ be the set of vertices along the string $C_{b}$ and $\{P_{c}\}$ be the set of plaquettes whose centers lie along $C_{b}$, where $c\in\{r,g\}$. Denote by $C^{v}_{rg}$ the set of red and green links which are incident on vertices along $C_{b}$ i.e., $C^{v}_{rg}=\{\ell_{c}:\ell_{c}\in{{}^{*}}v_{c}(c=r,g)\}$ for $v_{c}\in\{V_{c}\}$. Denote by $C^{p}_{rg}$ those red and green links which are shared by pairs of plaquettes along $C_{b}$ i.e., $C^{p}_{rg}=\{\ell_{c}:\ell_{c}\in\partial p_{c}\,\&\,\ell_{c}\in\partial p^{\prime}_{c}(c=r,g)\}$ for $p_{c}\neq p^{\prime}_{c}\in\{P_{c}\}$. Then $C_{rg}=C^{v}_{rg}\cap C^{p}_{rg}$.. The notation “$\ell_{r}>\ell_{g}$” indicates those pairs of red (connected to vertices $v_{r}$ along $C_{b}$) and green (connected to vertices $v_{g}$ along $C_{b}$) links in $C_{rg}$ for which $v_{r}$ is to the right of $v_{g}$. We can similarly define open-string operators that pair-create isolated red and green vertex defects.
@@
Using the definition of the plaquette stabilizers and the relation in Eq. (24), it is straightforward to verify that $\widetilde{X}_{u,v}$ creates defects only at the string’s endpoints: if $p$ is a plaquette that encloses a pair of edges in $C_{b}$, then $[A_{p_{r}},\widetilde{X}_{u,v}]=[A_{p_{g}},\widetilde{X}_{u,v}]=0$ and hence these stabilizers remain in their ground state after applying $\widetilde{X}_{u,v}$. Thus plaquette defects can occur only on the two plaquettes that enclose the vertices $u$ and $v$ (for which $\langle A_{p}\rangle=0$ after applying the $\widetilde{X}_{u,v}$ operator.)
@@
In other words, to create only a pair of vertex defects, a Pauli-X string acting on blue edges must be decorated with a sequence of $CZ$ gates that couple green and red edges along the string’s full length; the non-Abelian anyon string is thus inherently non-local. A bare undecorated string of $X$ operators, in contrast, cannot simply create a pair of point-like excitations; rather, it creates a veritable anyon soup (i.e. a linear combination of states with all different numbers of anyons) along the string’s length; the acyclic nature of the fusion rules in the $\mathcal{D}(D_{4})$ anyon theory is reflected in the fact that this soup contains only Abelian anyons. Note, however, that measuring the plaquette operators away from the string’s end-points has some probability to project all of these plaquettes onto their ground states, leaving only a pair of point-like defects at the string’s endpoints. In other words, these measurements, applied after the $X$ string, have some probability of effectively acting like the non-Abelian anyon string operator as shown in Fig. 7(b)666 Evidently, the intermediate plaquette operators commute with each other and can thus be measured simultaneously, as by assumption there are no blue vertex defects along the string.. Even for general measurement outcomes, re-applying the Pauli-$X$ string after measurements does not return the stabilizers to their initial states. Thus in a situation with active dynamics involving plaquette measurements, it is also not safe to assume that what began as a simple string of Pauli-$X$ errors acting on the ground state can be undone simply by re-applying this same string. This is the essence of the complications that we will encounter, relative to the Toric code, in constructing our decoder below.
@@
The full spectrum of point-like excitations (or superselection sectors) in this model can be described in terms of combinations of the three colors of vertex and plaquette defects, yielding the twenty-two anyons of the $D_{4}$ TO [84]. The correspondence between the the lattice defects discussed above and the conventional labels for these anyons by a conjugacy class and irreducible representation of its centralizer is summarized in Appendix B (see Table 1).
@@ IV.2 Error Correction Protocol
Figure 8: Correction protocol for $X$ errors in the $D_{4}$ model. (a) X-leaf move: A string of $X$ erasure errors, shown using thick blue lines, has vertex defects $B_{v}=-1$ at each endpoint. The highlighted plaquettes of red and green color along the length of the string are in a superposition state, as described by Eq. (25). A blue vertex $v_{b}$ has a single $X$ flag on the surrounding blue edges. This flag is removed and the vertex defect (if present) is displaced by applying $X$ to the corresponding qubit. Following this, new $Z$ flags are raised on edges $e_{0},e_{1},d_{0}$ and $d_{1}$ (wavy lines) to herald potential plaquette defects created during this correction step. (b) X-loop move. Left: a configuration with $X$ flags on both edges $(e_{0}$ and $e_{1})$ above and to the left of vertex $v$, which will result in an $X$-loop move. Here, we show a configuration that has a loop of blue $X$ flags, along with two possible $X$ error strings (colored in purple and yellow) consistent with the flags that can generate the pair of defects shown using blue triangles. Right: the loop move removes the flags on $e_{0}$ and $e_{1}$, and displaces any vertex defects up and to the right by applying $X_{e_{0}}$. Wavy lines indicate the edges $\Xi[v]=\{d_{0},d_{1},d_{2},d_{5},j_{1},j_{2}\}$ surrounding the vertex where new $Z$ flags are raised (if not already present). Additionally, as in Fig. 3, new $X$ flags on the outer edges of the plaquette are raised if they were not already present (in this case, a single new $X$ flag, highlighted in cyan).
@@
Figure 9: Correction protocol for $Z$ errors in the $D_{4}$ model. The dots indicate the plaquette defect $A_{p_{r}}=-1$ and wavy lines represent a $Z$ flag on the edge. The correction proceeds at a plaquette $p$ only if $X$ flags on all qubits (highlighted in green and blue) connected to the vertices $\mathcal{I}(p)$ enclosed by the plaquette $p$ are in the state $n^{f}_{X}=0$. (a) Z-leaf move: The defect on plaquette $p$ is displaced in the direction of the edge that has $Z$ flag on it. (b) Z-loop moves: Four different configurations where plaquette $p$ has $Z$ flags on edges $(e_{0},e_{1}),(e_{1},e_{2}),(e_{0},e_{2})$, and $(e_{0},e_{1},e_{2})$ are shown in sub-panels $(i)-(iv)$. The plaquette defect is pushed towards the topmost neighbor via application of $Z$. New $Z$ flags (shown in red) are raised on the edges $d_{0},d_{1}$ outside the plaquette, similarly to the Toric code protocol (see Fig.4).
@@
We now describe the local correction protocol that we use to stabilize a steady-state that lies in the same mixed-state phase as the ground space of the $D_{4}$ Hamiltonian Eq. (20), for noise rates below a finite threshold. Note that the heralded noise channels are the same as in the Toric code (Eqs. (2),(3)), aside from additional colour indices. Below, the colour index $c$ is suppressed for convenience wherever it does not appear explicitly.
@@
As with the correction protocol for the 2d Toric code (see Sec. III.2), our correction protocol for $D_{4}$ involves two types of processes. The first are leaf moves, which shrink strings of flags from their end-points inwards, while simultaneously moving any nearby stabilizer defects. The second are moves that can break open loops of flags, while pushing the corresponding stabilizer defects in a fixed direction (chosen to be the North-Western corner here) such that they necessarily remain adjacent to at least one flagged edge. In the $D_{4}$ case however, a new feature arises due the fact that the plaquette operators $A_{p}$ are “decorated.”
@@
We begin by discussing how the correction protocol acts on $X$ flags and the associated vertex defects. The basic idea underlying the correction protocol for $X$ flags and vertex defects is similar to that used in the Toric code. At a vertex where only one of the edges hosts an $X$ flag, we first measure $B_{v}$, then apply $X$ on the flagged edge if $B_{v}=-1$, and finally remove the flag – exactly as we did for the Toric code. The leaf operations must be supplemented with an $X$ loop move to ensure that the protocol can remove contractible loops of flags. As noted above, however, these correction steps may inadvertently create new $Z$-errors: thus, after this correction step we must also raise $Z$ flags to indicate the locations of the possible new $Z$-errors, as shown in Fig. 8.
We now describe these two types of moves in more detail. The loop moves act on the $X$ flags and $B_{v}$ defects exactly as for the Toric code. An example of the resulting loop moves, including the positions of the new $Z$ flags, is shown in Fig. 8 (b). There, a vertex $v_{b}$ on sublattice 1 (sublattice formed by south-east corners of blue honeycombs) of the blue honeycomb lattice has $X$ flags on the blue edges above it ($e_{0}$) and to its left ($e_{1}$) and no flag $X$ flag on the remaining edge $e_{2}$, ensuring that a loop move will be applied. The loop move removes the two $X$ flags on these edges, adds $X$ flags on the remaining four edges of the plaquette (if these are not already present), and pushes the vertex defect upwards and to the right by applying $X_{e_{0}}$. If the defect was created by an $X_{e_{0}}$ error, then we have exactly undone this error. However, if the defect was generated by an $X_{e_{1}}$ error, this correction leads to a net application of $X_{e_{0}}X_{e_{1}}$ to the initial state, in spite of the fact that the edges $e_{0}$ and $e_{1}$ no longer carry $X$ flags. From Eq. (25), we see that subsequent measurements of plaquette stabilizers (which occur during the $Z$ correction step described later) on this state will result in
$\begin{split}(1+s_{g}A_{p_{g}})X_{e_{0}}X_{e_{1}}|\psi_{0}\rangle&=(1+s_{g}Z_{j_{1}}Z_{j_{2}})X_{e_{0}}X_{e_{1}}|\psi_{0}\rangle\\ (1+s_{r}A_{p_{r}})X_{e_{0}}X_{e_{1}}|\psi_{0}\rangle&=(1+s_{r}Z_{d_{0}}Z_{d_{5}})X_{e_{0}}X_{e_{1}}|\psi_{0}\rangle\\ (1+s^{\prime}_{r}A_{p^{\prime}_{r}})X_{e_{0}}X_{e_{1}}|\psi_{0}\rangle&=(1+s^{\prime}_{r}Z_{d_{1}}Z_{d_{2}})X_{e_{0}}X_{e_{1}}|\psi_{0}\rangle,\end{split}$ (27)
@@
where $L^{\text{TC}(c)}$ is the Toric code jump operator defined in Eq. (8) and $\Xi(v_{c})$ represents the set of six qubits in Eq. (27), where new $Z$ flags are added if not already present. The bit string $\mathbf{m^{z}}\in\{0,1\}^{\otimes 6}$ accounts for all possible configurations of pre-existing $Z$ flags on the edges in $\Xi(v_{c})$ where we recall the relation $f^{+}_{z}f^{-}_{Z}=n^{f}_{Z}$. Similar to Toric code case, the loop moves are applied to vertices of only one of the sublattices of each $c$-colored honeycomb lattice.
The process for adding flags after a leaf move is similar. As noted before, the vertex defects are corrected either along the path they were generated or along an alternative path that is equivalent to the original one, up to a product of stabilizers. While the former correction leaves the state unchanged, the latter results in the action of closed loops of Pauli $X$ operators. Thus new $Z$ flags must also be added after leaf moves. Specifically, after a leaf move that results in overall application of $X_{j_{b}}$, measuring $A_{p_{g}}$ or $A_{p_{r}}$ on one of the plaquettes centered on the two endpoints of the edge $e_{j_{b}}$ gives:

### v2

@@ IV.1 $D_{4}$ Topological Order: Quasi-stabilizer model and anyons
Figure 6: Stabilizers for the $D_{4}$ model: qubits are located on edges of three interpenetrating honeycomb lattices, depicted in red, blue, and green. We use PBC in both vertical and horizontal directions, as defined in Eq. (2), for each of these lattices. A vertex operator $B_{v_{c}}$ on sublattice $1$ of each lattice color is shown in the upper left; the vertex operators on sublattice 22 are products of $Z$ operators on the remaining three edges at each vertex. A plaquette operator $A_{p_{c}}$ for each color is shown on the bottom left. Pauli $X$ operators act on the shaded $c$ colored edges; bold black lines indicate the two-qubit control-Z gates ($CZ_{j,k}=\frac{1}{2}(1+Z_{j}+Z_{k}-Z_{j}Z_{k})$) acting on pairs of qubits inside the plaquette. The box in the upper right shows plaquette operators that fail to commute, and instead satisfy Eq. (23) with $c=\text{green},c^{\prime}=\text{blue},c^{\prime\prime}=\text{red}$.
We begin with a brief overview of the properties of $D_{4}$ topological order, as well as the operators that can be used to stabilize $D_{4}$ topologically ordered ground states on the lattice in a conventional qubit architecture. $D_{4}$ TO describes the phase of matter obtained by considering a discrete gauge theory with a gauge group $D_{4}$ (i.e., the dihedral group with 8 elements) in $d=2$ spatial dimensions. The point-like excitations, or anyons, in this theory can be understood in terms of their (electric) charge and (magnetic) flux. The pure flux excitations are associated with the four non-trivial conjugacy classes of $D_{4}$, while pure charges are associated with the four non-trivial irreducible representations (irreps) of $D_{4}$. The remaining excitations, known as dyons, are labeled by a choice of conjugacy class together with a non-trivial representation of its centralizer. For $D_{4}$, this leads to a total of twenty-one non-trivial anyons (or twenty-two total anyons, including the identity). A more comprehensive description of this anyon theory is provided in Appendix B.
Since $D_{4}$ is a non-Abelian group, some of these anyons are non-Abelian: when a pair of such anyons is brought together (or fused), the resulting anyon type is not uniquely determined. For example, $D_{4}$ has one 2-dimensional irrep; two charges of this type can be combined in multiple ways, yielding different charge types (i.e., multiple irreps), in a manner analogous to the addition of angular momentum for irreps of SU(2).
@@
In the remainder of this Section, we review a specific model Hamiltonian, introduced in Ref. [107], that realizes the $\mathcal{D}(D_{4})$ topological phase. This model has two advantages: first, the local Hilbert space is composed of qubits, rendering it simpler to realize on current quantum hardware [67]; and second, the Hamiltonian can be written as a sum of local “check” operators, each of which individually take eigenvalue $1$ in the ground state subspace despite not commuting in the full Hilbert space.
The model is defined on three inter-penetrating honeycomb lattices with a qubit placed on each edge, as shown in Fig. 6. For convenience, the three honeycomb lattices are shown in different colors (red, green, and blue in the Figure). With each color, we associate two types of check operators that stabilize the ground states. First, the vertex check centered at vertex $v_{c}$ is defined as
@@
Note that the plaquette check operators defined above are non-Pauli “quasi-stabilizers”. Specifically, unlike standard stabilizer operators on an $N$ qubit Hilbert space, these check operators are not elements of the $N$ qubit Pauli group, and do not commute with each other as operators. However, they do commute within the ground-state subspace (see below for details), and stabilize the $D_{4}$ ground states. Thus, in the interest of brevity (and in an abuse of notation), we will often refer to $A_{p}$ and $B_{v}$ operators simply as stabilizers and will emphasize the non-trivial consequences of their “quasi-stabilizer” nature whenever necessary.
@@
$B_{v_{c}},B_{v^{\prime}_{c^{\prime}}}]=[B_{v_{c}},A_{p_{c^{\prime}}}]=0\,\,\forall\,p,v,v^{\prime},c,c^{\prime}\,.$ (22)
@@
[…]
where $A_{p_{r}}$ is a red plaquette that encloses the blue edge $j_{b}$ and $k_{g},k^{\prime}_{g}\in\ ^{*}\mathcal{I}(p_{r})$ denote the green edges inside this red plaquette that are connected by $CZ$ gates to the edge $j_{b}$. Thus, in configurations with $Z_{k_{g}}Z_{k^{\prime}_{g}}=+1$, measuring $A_{p_{r}}$ after the $X_{j_{b}}$ error acts on the ground state $|\psi_{D_{4}}\rangle$ will find no plaquette error, while in configurations with $Z_{k_{g}}Z_{k^{\prime}_{g}}=-1$ a plaquette error will be found. Since both configurations occur in the ground state, $X_{j_{b}}$ leaves this red plaquette in a superposition of having and not having a plaquette defect. Similarly, acting with a longer string of $X$ operators on blue edges will result in a pair of vertex violations at the string’s endpoints, together with a superposition of violated and unviolated plaquette stabilizers along the string’s length, as depicted by the highlighted plaquettes in Fig. 8(a).
@@
Here, $C_{b}$ is an open string along the blue honeycomb lattice that connects vertices $u$ and $v$, and $\ell_{b}\in C_{b}$ denotes links along this path (see Fig. 7(a)). The set $C_{rg}$ consists of those links on the red and green lattices which are incident on vertices along $C_{b}$ and are also either enclosed by or are on the boundary of two plaquettes that enclose vertices in $C_{b}$ [^6] Let $\{V_{c}\}$ be the set of vertices along the string $C_{b}$ and $\{P_{c}\}$ be the set of plaquettes whose centers lie along $C_{b}$, where $c\in\{r,g\}$. Denote by $C^{v}_{rg}$ the set of red and green links which are incident on vertices along $C_{b}$ i.e., $C^{v}_{rg}=\{\ell_{c}:\ell_{c}\in{{}^{*}}v_{c}(c=r,g)\}$ for $v_{c}\in\{V_{c}\}$. Denote by $C^{p}_{rg}$ those red and green links which are shared by pairs of plaquettes along $C_{b}$ i.e., $C^{p}_{rg}=\{\ell_{c}:\ell_{c}\in\partial p_{c}\,\&\,\ell_{c}\in\partial p^{\prime}_{c}(c=r,g)\}$ for $p_{c}\neq p^{\prime}_{c}\in\{P_{c}\}$. Then $C_{rg}=C^{v}_{rg}\cap C^{p}_{rg}$.. The notation $\ell_{r}>\ell_{g}$” indicates those pairs of red (connected to vertices $v_{r}$ along $C_{b}$) and green (connected to vertices $v_{g}$ along $C_{b}$) links in $C_{rg}$ for which $v_{r}$ is to the right of $v_{g}$. We can similarly define open-string operators that pair-create isolated red and green vertex defects.
@@
Using the definition of the plaquette stabilizers and the relation in Eq. (25), it is straightforward to verify that $\widetilde{X}_{u,v}$ creates defects only at the string’s endpoints: if $p$ is a plaquette that encloses a pair of edges in $C_{b}$, then $A_{p_{r}},\widetilde{X}_{u,v}]=[A_{p_{g}},\widetilde{X}_{u,v}]=0$ and hence these stabilizers remain in their ground state after applying $\widetilde{X}_{u,v}$. Thus, plaquette defects can occur only on the two plaquettes that enclose the vertices $u$ and $v$ (for which $\langle A_{p}\rangle=0$ after applying the $\widetilde{X}_{u,v}$ operator.)
@@
In other words, to create only a pair of vertex defects, a Pauli-X string acting on blue edges must be decorated with a sequence of $CZ$ gates that couple green and red edges along the string’s full length; the non-Abelian anyon string is thus inherently non-local. A bare undecorated string of $X$ operators, in contrast, cannot simply create a pair of point-like excitations; rather, it creates a veritable anyon soup (i.e. a linear combination of states with all different numbers of anyons) along the string’s length; the acyclic nature of the fusion rules in the $\mathcal{D}(D_{4})$ anyon theory is reflected in the fact that this soup contains only Abelian anyons. Note, however, that measuring the plaquette operators away from the string’s end-points has some probability to project all of these plaquettes onto their ground states, leaving only a pair of point-like defects at the string’s endpoints. In other words, these measurements, applied after the $X$ string, have some probability of effectively acting like the non-Abelian anyon string operator as shown in Fig. 7(b) [^7] Evidently, the intermediate plaquette operators commute with each other and can thus be measured simultaneously, as by assumption there are no blue vertex defects along the string.. Even for general measurement outcomes, re-applying the Pauli-$X$ string after measurements does not return the stabilizers to their initial states. Thus in a situation with active dynamics involving plaquette measurements, it is also not safe to assume that what began as a simple string of Pauli-$X$ errors acting on the ground state can be undone simply by re-applying this same string. This is the essence of the complications that we will encounter, relative to the Toric code, in constructing our decoder below.
@@
The full spectrum of point-like excitations (or superselection sectors) in this model can be described in terms of combinations of the three colors of vertex and plaquette defects, yielding the twenty-two anyons of the $D_{4}$ TO [88]. The correspondence between the lattice defects discussed above and the conventional labels for these anyons by a conjugacy class and irreducible representation of its centralizer is summarized in Appendix B (see Table 1).
@@ IV.2 Error Correction Protocol
Figure 8: Correction protocol for $X$ errors in the $D_{4}$ model. (a) X-leaf move: A string of $X$ erasure errors, shown using thick blue lines, has vertex defects $B_{v}=-1$ at each endpoint. The highlighted plaquettes of red and green color along the length of the string are in a superposition state, as described by Eq. (26). A blue vertex $v_{b}$ has a single $X$ flag on the surrounding blue edges. This flag is removed, and the vertex defect (if present) is displaced by applying $X$ to the corresponding qubit. Following this, new $Z$ flags are raised on edges $e_{0},e_{1},d_{0}$ and $d_{1}$ (wavy lines) to herald potential plaquette defects created during this correction step. (b) X-loop move. Left: a configuration with $X$ flags on both edges $(e_{0}$ and $e_{1})$ above and to the left of vertex $v$, which will result in an $X$-loop move. Here, we show a configuration that has a loop of blue $X$ flags, along with two possible $X$ error strings (colored in purple and yellow) consistent with the flags that can generate the pair of defects shown using blue triangles. Right: the loop move removes the flags on $e_{0}$ and $e_{1}$, and displaces any vertex defects up and to the right by applying $X_{e_{0}}$. Wavy lines indicate the edges $\Xi[v]=\{d_{0},d_{1},d_{2},d_{5},j_{1},j_{2}\}$ surrounding the vertex where new $Z$ flags are raised (if not already present). Additionally, as in Fig. 3, new $X$ flags on the outer edges of the plaquette are raised if they were not already present (in this case, a single new $X$ flag, highlighted in cyan).
@@
Figure 9: Correction protocol for $Z$ errors in the $D_{4}$ model. The dots indicate the plaquette defect $A_{p_{r}}=-1$, and wavy lines represent a $Z$ flag on the edge. The correction proceeds at a plaquette $p$ only if $X$ flags on all qubits (highlighted in green and blue) connected to the vertices $\mathcal{I}(p)$ enclosed by the plaquette $p$ are in the state $n^{f}_{X}=0$. (a) Z-leaf move: The defect on plaquette $p$ is displaced in the direction of the edge that has $Z$ flag on it. (b) Z-loop moves: Four different configurations where plaquette $p$ has $Z$ flags on edges $(e_{0},e_{1}),(e_{1},e_{2}),(e_{0},e_{2})$, and $(e_{0},e_{1},e_{2})$ are shown in sub-panels (i)(iv)(i)-(iv). The plaquette defect is pushed towards the topmost neighbor via application of $Z$. New $Z$ flags (shown in red) are raised on the edges $d_{0},d_{1}$ outside the plaquette, similarly to the Toric code protocol (see Fig.4).
@@
We now describe the local correction protocol that we use to stabilize a steady-state that lies in the same mixed-state phase as the ground space of the $D_{4}$ Hamiltonian Eq. (21), for noise rates below a finite threshold. Note that the heralded noise channels are the same as in the Toric code (Eqs. (5),(6)), aside from additional color indices. Below, the color index $c$ is suppressed for convenience wherever it does not appear explicitly.
@@
As with the correction protocol for the 2d Toric code (see Sec. III.2), our correction protocol for $D_{4}$ involves two types of processes. The first are leaf moves, which shrink strings of flags from their end-points inwards, while simultaneously moving any nearby stabilizer defects. The second are moves that can break open loops of flags, while pushing the corresponding stabilizer defects in a fixed direction (chosen to be the North-Western corner here) such that they necessarily remain adjacent to at least one flagged edge. In the $D_{4}$ case, however, a new feature arises due to the fact that the plaquette operators $A_{p}$ are “decorated.”
@@
We begin by discussing how the correction protocol acts on $X$ flags and the associated vertex defects. The basic idea underlying the correction protocol for $X$ flags and vertex defects is similar to that used in the Toric code. At a vertex where only one of the edges hosts an $X$ flag, we first measure $B_{v}$, then apply $X$ on the flagged edge if $B_{v}=-1$, and finally remove the flag – exactly as we did for the Toric code. The leaf operations must be supplemented with an $X$ loop move to ensure that the protocol can remove contractible loops of flags. As noted above, however, these correction steps may inadvertently create new $Z$-errors: thus, after this correction step, we must also raise $Z$ flags to indicate the locations of the possible new $Z$-errors, as shown in Fig. 8.
We now describe these two types of moves in more detail. The loop moves act on the $X$ flags and $B_{v}$ defects exactly as for the Toric code. An example of the resulting loop moves, including the positions of the new $Z$ flags, is shown in Fig. 8 (b). There, a vertex $v_{b}$ on sublattice 1 (sublattice formed by south-east corners of blue honeycombs) of the blue honeycomb lattice has $X$ flags on the blue edges above it ($e_{0}$) and to its left ($e_{1}$) and no $X$ flag on the remaining edge $e_{2}$, ensuring that a loop move will be applied. The loop move removes the two $X$ flags on these edges, adds $X$ flags on the remaining four edges of the plaquette (if these are not already present), and pushes the vertex defect upwards and to the right by applying $X_{e_{0}}$. If the defect was created by an $X_{e_{0}}$ error, then we have exactly undone this error. However, if the defect was generated by an $X_{e_{1}}$ error, this correction leads to a net application of $X_{e_{0}}X_{e_{1}}$ to the initial state, in spite of the fact that the edges $e_{0}$ and $e_{1}$ no longer carry $X$ flags. From Eq. (26), we see that subsequent measurements of plaquette stabilizers (which occur during the $Z$ correction step described later) on this state will result in
$(1+s_{g}A_{p_{g}})X_{e_{0}}X_{e_{1}}|\psi_{0}\rangle$ $=(1+s_{g}Z_{j_{1}}Z_{j_{2}})X_{e_{0}}X_{e_{1}}|\psi_{0}\rangle$ (28)
$(1+s_{r}A_{p_{r}})X_{e_{0}}X_{e_{1}}|\psi_{0}\rangle$ $=(1+s_{r}Z_{d_{0}}Z_{d_{5}})X_{e_{0}}X_{e_{1}}|\psi_{0}\rangle$
$(1+s^{\prime}_{r}A_{p^{\prime}_{r}})X_{e_{0}}X_{e_{1}}|\psi_{0}\rangle$ $=(1+s^{\prime}_{r}Z_{d_{1}}Z_{d_{2}})X_{e_{0}}X_{e_{1}}|\psi_{0}\rangle,$
@@
where $L^{\text{TC}(c)}$ is the Toric code jump operator defined in Eq. (9) and $\Xi(v_{c})$ represents the set of six qubits in Eq. (28), where new $Z$ flags are added if not already present. The bit string $\mathbf{m^{z}}\in\{0,1\}^{\otimes 6}$ accounts for all possible configurations of pre-existing $Z$ flags on the edges in $\Xi(v_{c})$ where we recall the relation $f^{+}_{z}f^{-}_{Z}=n^{f}_{Z}$. Similar to the Toric code case, the loop moves are applied to the vertices of only one of the sublattices of each $c$-colored honeycomb lattice.
The process for adding flags after a leaf move is similar. As noted before, the vertex defects are corrected either along the path they were generated or along an alternative path that is equivalent to the original one, up to a product of stabilizers. While the former correction leaves the state unchanged, the latter results in the action of closed loops of Pauli $X$ operators. Thus, new $Z$ flags must also be added after leaf moves. Specifically, after a leaf move that results in overall application of $X_{j_{b}}$, measuring $A_{p_{g}}$ or $A_{p_{r}}$ on one of the plaquettes centered on the two endpoints of the edge $e_{j_{b}}$ gives: